# Stringlet Excitation Model of the Boson Peak

### v2

@@ II The stringlet model
We notice that a similar expression for the BP frequency as in Eqs.(9)-(10) was suggested before by Granato Granato (1992), Hong et al. Hong, Novikov, and Sokolov (2011a, b) and Kalampounias et al. Kalampounias, Yannopoulos, and Papatheodorou (2006a) on a more heuristic basis (See also Refs.Malinovsky et al. (1988); Malinovsky and Sokolov (1986) and Ref.Martin and Brenig (1974) for the original idea of introducing a length scale associated with the boson peak.). The length-scale appearing in Eqs.(9)-(10) has been interpreted by Hong et al.Hong, Novikov, and Sokolov (2011a, b) in recent work as a “cooperativity length scale” $\xi$ associated with cooperative motion associated with the growing activation energy of the alpha relaxation time of glass-forming liquids postulated in the Adam-Gibbs model Adam and Gibbs (1965). However, there is some evidence that this characteristic length defined from an equation of the assumed form of Eqs.(9) and (10) increases with temperature, as observed for the stringlet length $\lambda$ in simulations Zhang et al. (2021a); Zhang and Douglas (2013b). In contrast, the activation and the associated scale of collective motion in glass-forming liquids grows upon cooling, Refs.Betancourt P., Starr, and Douglas (2018); Pazmiño Betancourt, Douglas, and Starr (2014), a trend that is contrary to the intuitively attractive idea of identifying $\xi$ with the size of the “cooperatively rearranging regions” of Adam-Gibbs. Independent of these experimental observations on the boson peak, string-like collective motion whose extent grows with temperature has been inferred from low angle inelastic coherent neutron scattering measurements Russina et al. (2000). Notice that the numerical prefactor in Eqs.(9)-(10), defining the characteristic length associated with the boson peak, is completely unspecified in previous studies, while it is specified exactly in our theory. We finally remark that our theoretical approach presents two important improvements with respect to the earlier phenomenological studies reported above. First, the length scale appearing in the expression for the BP frequency is immediately identified from the stringlet length obtained from simulations observations. At the same time, the numerical prefactor describing the stringlet size distribution can be directly extracted from the simulation data. Compared to the previous literature, our theory is based on more physical arguments that go beyond a pure dimensional analysis or vague ideas about cooperative motion of some unspecified kind. Nevertheless, it is rather remarkable that the brilliant raw intuition behind the qualitative picture of Sokolov and coworkers Hong, Novikov, and Sokolov (2011a, b) anticipated the functional form predicted by our model.
@@
In summary, this simple model exhibits three striking predictions. (I) It gives an analytical and simple estimate for the BP frequency in both 2D and 3D systems, Eqs.(9)-(10). These results imply that the BP in a 3D system should appear at lower energies compared to a 2D system with the same length distribution and sound speed. (II) If we reasonably assume that the speed of the stringlet excitation is the same of that of transverse phonons, this model immediately predicts that $\omega_{BP}\sim G^{1/2}$ in the low temperature glass state where the stringlet average length (and the configurational entropy as well) are expected to become independent of temperature Betancourt et al. (2015); Betancourt P., Starr, and Douglas (2018); Douglas et al. (2016). However, the stringlet model also predicts that this simple scaling becomes progressively modified at temperatures approaching $T_{g}$, where the temperature dependence of $\lambda$ can no longer be neglected. Importantly, identifying the velocity of propagation of stringlet excitations with the average speed of transverse sound is a working hypothesis in our coarse-grained model. Nevertheless, this hypothesis can be motivated by previous observations in simulations Shintani and Tanaka (2008); Hu and Tanaka (2023) and measurements Sokolov et al. (1992, 1997) that the boson peak strongly correlates with the transverse sound modes. Moreover, we remark that Sokolov and coworkers Hong, Novikov, and Sokolov (2011a, b) have recently analyzed the boson peak in diverse materials with a view of defining a dynamical correlation length for collective motion based on an assumed expression in which the boson peak is taken to be proportional to the transverse sound velocity divided by a cooperative length scale (the analog of the stringlet “size” in our theory). (III) At highly elevated temperatures where we may expect the glass to become a liquid, however, the stringlet theory interestingly predicts a linear frequency dependence of the VDOS, and thus leads to a $1/\omega$ divergence of the reduced density of states, the density of states normalized by the Debye density of states. This scaling is consistent with recent modeling of the density of states so that the stringlet model appears to recover a physically correct behavior in the high temperature simple liquid as well as in the low temperature glass states.
@@
We next consider the capacity of the revised string model of the boson peak to quantitatively account for the frequency dependence of the boson peak. First, we obtain the value of the parameter $\lambda$ in Eq.(5) by fitting the distribution of stringlet length obtained from simulations. Then, we take the value for the velocity of transverse phonons $v_{T}$ which is obtained by fitting the dispersion relation obtained from simulations. Once these two parameters are known, and fixed by the simulations, our theory provides a parameter-free prediction for the BP frequency in both 2D and 3D amorphous solids, Eqs.(9)-(10). We can then directly compare the predicted value for $\omega_{BP}$, which we will denote as $\omega_{BP}^{th}$, with the experimental numbers obtained from the simulation data, which we will label as $\omega_{BP}^{exp}$.
We start by considering the simulation data for 2D and 3D zero temperature glasses presented in Hu and Tanaka (2022, 2023). In Fig.2, we present two explicit examples performed using the data for the 2D power law model (2DPL) in Hu and Tanaka (2022) and the 3D inverse power law model (3DIPL) in Hu and Tanaka (2023). By plausibly assuming that the average velocity of the stringlet propagation coincides with the speed of transverse sound, we obtain a prediction for the BP of the 2D glass given by $\omega_{BP}^{th}\approx 1.344$ which has to be compared with the reported simulation value $\omega_{BP}^{exp}=1.41$. The difference between the two is less than $5\,\%$.
For the 3D system, the agreement is even superior, showing $\approx 1\,\%$ uncertainty. The predictions of the theory and the compared experimental values are reported in Table 1.
Table 1: The theoretical prediction versus the experimental value for the BP frequency using the exponential fit, Eq.(5). The uncertainty $\Upsilon$ is defined as $\Upsilon=2|\omega_{BP}^{th}-\omega_{BP}^{exp}|/(\omega_{BP}^{th}+\omega_{BP}^{exp})$. All quantities presented here are in reduced Lennard-Jones units.
$\lambda$ $v_{T}$ $\omega_{BP}^{th}$ $\omega_{BP}^{exp}$ $\Upsilon$
3DIPL Hu and Tanaka (2023) $1.335$ $4.23$ $2.487$ $2.46$ $1\,\%$
2DPL Hu and Tanaka (2022) $3.717$ $4.78$ $1.344$ $1.41$ $4.7\,\%$
The estimates from the theoretical model are in remarkable agreement, i.e., agreement to within $5\,\%$ uncertainty, with the value for the BP frequency obtained from simulations. It is emphasized that, once the distribution of length is extracted from the experimental data and the value of the speed of transverse sound is taken from the simulation dispersion relation, the theory has no free parameters.
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
In order to test further our theory, we have extracted the simulation data from Hu and Tanaka (2022, 2023) and compared them directly to the total VDOS with our predictions. In order to do so, we added a Debye contribution to the total DOS:
$g_{\text{Debye}}(\omega)=\mathcal{A}\,\omega^{2},$ (17)
where the Debye level $\mathcal{A}$ is fixed using the numerical data, and a constant in frequency renormalization factor in front of the stringlet contribution as well. Additionally, as evident from the figures in the bottom panel of Fig.2, a more careful analysis reveals that the fit to an exponential becomes rather uncertain for longer strings which are relatively rare in the simulation sampling, i.e., $l>15$ in reduced LJ units. This was already observed in Pazmiño Betancourt, Douglas, and Starr (2014), where a combination of a power-law with an exponential was considered and shown to fit the data better. In order to investigate this point further, we have performed a second analysis in which the stringlet length distribution $p(l)$ is not fitted with the exponential form, Eq.(5), but rather with a power-law corrected expression as in Eq.(11).
We compare our analytic expression for the full reduced density of sates to our analytic model prediction in the top panels of Fig.2. The simulation data are shown with black symbols, the theoretical prediction based on the exponential distribution, Eq.(5), is presented with solid black lines, the theoretical prediction based on the power-law corrected distribution Eq.(11) by the dashed black lines, and finally the gray vertical bar indicates the position of the BP. In the two-dimensional model (2DPL) the comparison between the simulation data and our theoretical prediction is rather good, specially for the exponential fit. The fitting corrected by a power-law term seems to give a slightly better result for frequencies below the BP (corresponding to large $l$ in the distribution function $p(l)$), but a worse result for frequencies around and above the BP. In the three dimensional system, the comparison between the theory and the simulations is good below $\omega\approx 4$ (in reduced units), but it becomes poor for large frequencies, much above the BP scale. In particular, the simulation data seem to decay much faster than our theoretical expectations. This discrepancy can be easily understood by noticing that our theoretical prediction does not include any UV cutoff to take into account the microscopic scale of the system. By adding a phenomenological Debye cutoff to the total density of states, $\exp\left(-\omega^{2}/\omega_{D}^{2}\right)$, this discrepancy immediately vanishes as shown by the red dashed curve in the top right panel of Fig.2. Finally, we notice that this discrepancy does not appear in the 2D model because the comparison does not extend to large enough frequencies.
@@ III Damped dynamics in cooled liquids above the glass transition
Inspired by these positive outcomes in the comparisons of the stringlet model to simulations deep in the glass state, we next consider simulations for a simulated Al-Sm metallic glass-forming material in a temperature range above $T_{g}$ discussed previously in Zhang et al. (2021a) (see also Zhang et al. (2021b)). By using the experimental values reported in Table 2, see also Fig.3, we are again able to provide a parameter-free theoretical prediction for the BP frequency in the limit of zero damping, which is shown in the last column of the same table.
Table 2: The parameters for the Al-Sm system at finite temperature Zhang et al. (2021a). $\lambda$ is reported in nanometers, the velocity in m/s and the frequencies in meV. The last column indicates the theoretical prediction for zero stringlet damping which has to be compared with the previous column reporting the corresponding simulation values.
Al-Sm Zhang et al. (2021a) $\lambda$ $v_{T}$ $\omega_{BP}^{exp}$ $\omega_{BP}^{th}$ $(\gamma=0)$
@@
The stringlet length distribution and the average stringlet length $\lambda$ are shown for different temperatures in the left panel of Fig.3. Moreover, we obtained the values for the density $\rho$ and the shear modulus $G$ from Zhang et al. (2021a, b) and estimated the transverse speed of sound using the well-known relation:
@@
Refer to captionRefer to caption
Figure 3: Left panel: The distribution of the stringlet lengths for the finite temperature Al-Sm system in Zhang et al. (2021a) and the fits to the exponential form, Eq.(5). The inset shows the dependence of the average stringlet length $\lambda$ as a function of temperature. Right panel: The speed of transverse sound, extracted from the simulation data using Eq.(18), as a function of temperature. The values of $\lambda$ and $v_{T}$ are reported explicitly in Table 2.
Refer to captionRefer to caption
Figure 4: Left panel: The phenomenological damping parameter used in the theoretical predictions at high temperature, above $T_{g}$, for the Al-Sm system Zhang et al. (2021a). The colors match those used in the bottom panel. Right panel: The comparison between the theoretical predictions and the simulation results for the finite temperature Al-Sm system in Zhang et al. (2021a). The black symbols are the simulation data. The orange line indicates the theoretical results in absence of stringlet damping, $\gamma=0$. The pink and red line indicate the theoretical results in presence of damping, as explained in the text.
@@
One qualitative finding is evident from Table 2 (see also Fig.4). Simulation observations cannot be described by a theory neglecting stringlet damping. The theoretical predictions neglecting damping overestimate the observed values of $\omega_{BP}$ by more than a factor of two. The reason for this deviation is easy to understand. As opposed to the simulations of Hu and Tanaka Hu and Tanaka (2022, 2023), which were performed in the limit of $T=0$ where stringlet damping can be reasonably neglected, the simulations for the Al-Sm metallic glass forming material Zhang et al. (2021a) were performed at relatively high temperatures corresponding to a highly cooled glass-forming metallic glass liquid Zhang et al. (2021a). At such temperatures, it is unreasonable to assume $\gamma=0$ even as a rough approximation. The treatment of the boson peak at higher temperatures is complicated by the fact that the stringlet length, the transverse sound velocity and the extent of damping described by $\gamma$ can all be expected to vary appreciably with temperature (see Figs. 3 and 4 for the temperature variation observed in the Al-Sm metallic glass-forming liquid). Nevertheless, as we will see below, we can still gain important physical insights into the variation of the density of states and the boson peak with temperature by considering the stringlet model with finite damping $\gamma$ and by considering a simplified description for it.
Before continuing, let us emphasize that our analysis does not take into account the existence of hydrodynamically-defined “defects”, such as the often discussed four-leaf quadrupole defects Lerner and Bouchbinder (2021). Nevertheless, as argued in Ref.Hu and Tanaka (2022), these relatively large scale structures apparently correspond to a frequency range much below the boson peak, and thus not germane to the present analysis. In particular, we do not expect this type of defects to have a significant impact in the frequency range near where the boson peak is exhibited in the reduced density of states. It is nevertheless plausible that these additional defects might contribute to the stringlet damping, especially at elevated temperatures and/or large frequencies, and it is also highly likely that these defects are highly significant in connection with understanding plastic deformation in glassy materials. We certainly do not mean to imply these defects are not important in glass-forming materials. Finally, we note that the existence of quadrupole defects defined in relation to elastic distortions in the continuum field describing the material provides a well defined example of how defects might arise without any crystalline structure necessarily existing in the material in order to define a defect. This would seem to imply that a similar non-crystalline based definition might apply to stringlets so that these structures might genuinely be termed “defects”. This possibility remains to be established.
@@
$\frac{g_{s}(\omega)}{\omega^{2}}=\frac{p(0)\,\pi\,(d-1)\,v}{\gamma^{2}}\frac{1}{\omega}+\dots$ (19)
@@
Note also that Eq.(19) implies that the VDOS is linear in frequency at low energy in the high temperature liquid state so that our stringlet model exhibits the correct physical behavior in both the low temperature solid and the high temperature liquid states. This behavior is consistent with the experimental observations in classical liquids Phillips et al. (1989); Stamper et al. (2022); Jin et al. (2023) where it has been rationalized based on the concept of unstable modes Keyes (1997); Zaccone and Baggioli (2021). Interestingly, string-like excitations have been also discussed in glass-forming liquids at low temperatures Donati et al. (1998); Miyagawa et al. (1988); Pazmiño Betancourt, Douglas, and Starr (2014), so that it would be important to study further the BP in relation to stringlet dynamics by heating such cooled liquids to watch how the boson peak evolves in shape and eventually disappears. Notably, this disappearance of the boson peak upon heating is well-established experimentally Hong, Novikov, and Sokolov (2011a); Rufflé et al. (2010); Chumakov et al. (2004); Zanatta et al. (2011); Kalampounias, Yannopoulos, and Papatheodorou (2006a); González-Jiménez et al. (2023). Similarities to the occurrence of glassy anomalies in materials undergoing commensurate-incommensurate transitions Cano and Levanyuk (2004); Reményi et al. (2015); Etrillard et al. (1996); Jiang et al. (2023); Biljaković et al. (2012) and we suggest that the pursuit of this “analogy” might offer potential insights into the boson peak in glass-forming liquids. We next tentatively explore the extension to higher temperatures where we may expect damping of the stringlets.
@@
We start our discussion about the predictions of our stringlet model at elevated temperatures by noting that an increase of the average stringlet length $\lambda$ with temperature is not sufficient to reproduce the complete phenomenology of the “softening” of the BP, and the simultaneous emergence of a linear in $\omega$ VDOS as appropriate for liquids Stamper et al. (2022). At the outset, we admit that the observed softening of the BP upon heating might involve multiple mechanisms, among which the growth of the average stringlet length $\lambda$, the decrease of the velocity $v$, or the increase of the stringlet damping $\gamma$. From the simulation data, only the variation of $\lambda$ and $v$ with $T$ is known, but we cannot currently directly access information about the dependence of $\gamma$ on $T$. Thus, we cannot fully resolve how each mechanism contributes to the BP within the frame of our model without introducing a model of $\gamma(T)$. We introduce a tentative model of $\gamma(T)$ below to address this problem.
Refer to captionRefer to caption
@@
Figure 5: Left panel: The stringlet vibrational density of states normalized by Debye law $\omega^{2}$ from Eq.(6). Here, we have fixed $p(0)=v=\lambda=1$ and we move the damping $\gamma\in[10^{-4},0.25]$ from black to red. The arrows indicate the location of the BP and of the low-frequency $1/\omega$ divergence discussed in Eq.(19). Right panel: The apparent power law emerging at low frequency in presence of damping. In this figure, we have fixed $\lambda=v=p(0)=1$ and dialed the damping $\gamma\in[0,1]$. A Debye term $g_{Debye}(\omega)=\omega^{2}$ has been added on top of the stringlet contribution, Eq.(6).

### v3

@@ II The stringlet model
We notice that a similar expression for the BP frequency as in Eqs.(11)-(12) was suggested before by Granato [71], Hong et al. [60; 61] and Kalampounias et al. [72] on a more heuristic basis (See also Refs.[73; 74] and Ref.[75] for the original idea of introducing a length scale associated with the boson peak.). The length-scale appearing in Eqs.(11)-(12) has been interpreted by Hong et al.[60; 61] in recent work as a “cooperativity length scale” $\xi$ associated with cooperative motion associated with the growing activation energy of the alpha relaxation time of glass-forming liquids postulated in the Adam-Gibbs model [76]. However, there is some evidence that this characteristic length defined from an equation of the assumed form of Eqs.(11) and (12) increases with temperature, as observed for the stringlet length in simulations [32; 34]. In contrast, the activation and the associated scale of collective motion in glass-forming liquids grows upon cooling, Refs.[37; 59], a trend that is contrary to the intuitively attractive idea of identifying $\xi$ with the size of the “cooperatively rearranging regions” of Adam-Gibbs. Independent of these experimental observations on the boson peak, string-like collective motion whose extent grows with temperature has been inferred from low angle inelastic coherent neutron scattering measurements [77]. Notice that the numerical prefactor in Eqs.(11)-(12), defining the characteristic length associated with the boson peak, is completely unspecified in previous studies, while it is specified exactly in our model. We finally remark that our theoretical approach presents two important improvements with respect to the earlier phenomenological studies reported above. First, the length scale appearing in the expression for the BP frequency is immediately identified with the stringlet length obtained from simulations observations. At the same time, the numerical prefactor describing the stringlet size distribution can be directly extracted from the simulation data. Compared to the previous literature, our theoretical model is based on more physical arguments that go beyond a pure dimensional analysis or vague ideas about cooperative motion of some unspecified kind. Nevertheless, it is rather remarkable that the brilliant raw intuition behind the qualitative picture of Sokolov and coworkers [60; 61] anticipated the functional form predicted by our model.
@@
In summary, this simple model exhibits three striking predictions. (I) It gives an analytical and simple estimate for the BP frequency in both 2D and 3D systems, Eqs.(11)-(12). These results imply that the BP in a 3D system should appear at lower energies compared to a 2D system with the same length distribution and sound speed. (II) If we reasonably assume that the speed of the stringlet excitation is the same of that of transverse phonons, this model immediately predicts that $\omega_{BP}\sim G^{1/2}$ in the low temperature glass state where the stringlet average length (and the configurational entropy as well) are expected to become independent of temperature [38; 37; 39]. However, the stringlet model also predicts that this simple scaling becomes progressively modified at temperatures approaching $T_{g}$, where the temperature dependence of $\lambda$ can no longer be neglected. Importantly, identifying the velocity of propagation of stringlet excitations with the average speed of transverse sound is a working hypothesis in our coarse-grained model. Nevertheless, this hypothesis can be motivated by previous observations in simulations [29; 36] and measurements [78; 79] that the boson peak strongly correlates with the transverse sound modes. Moreover, we remark that Sokolov and coworkers [60; 61] have recently analyzed the boson peak in diverse materials with a view of defining a dynamical correlation length for collective motion based on an assumed expression in which the boson peak is taken to be proportional to the transverse sound velocity divided by a cooperative length scale (the analog of the stringlet “size” in our model). (III) At highly elevated temperatures where we may expect the glass to become a liquid, however, the stringlet model interestingly predicts a linear frequency dependence of the VDOS, and thus leads to a $1/\omega$ divergence of the reduced density of states, the density of states normalized by the Debye density of states. This scaling is consistent with recent modeling of the density of states so that the stringlet model appears to recover a physically correct behavior in the high temperature simple liquid as well as in the low temperature glass states.
Before proceeding with our results, we notice that the identification and characterization of the stringlets from simulations can be done in different ways and can be achieved from the analysis of the collective dynamics on the $\beta$ relaxation time scale. This is how the stringlets were originally defined and characterized both in liquids and amorphous solids, see for example Ref. [37] (in particular figure 3 from which the stringlet length can be extracted) and Ref. [80] for the original method and idea. In this sense, the stringlet distribution $p(l)$ is an independent input that can be obtained from simulations (without diagonalization of the dynamical matrix) and then used to predict the BP location in the density of states. We note that the diagonalization of the dynamical matrix involves a harmonic approximation which is best to avoid if possible. Tanaka and collaborators [35], more recently, introduced a new method to identify and characterize these stringlets by looking at the particle-level vibrational density of states. Regarding this second method, the identification of the stringlets with the boson peak necessitates a knowledge of the density of states and anomalous excess modes from those expected from Debye theory. There are a variety of methods of determining the density of states, however.
@@
We next consider the capacity of the revised string model of the boson peak to quantitatively account for the frequency dependence of the boson peak. First, we obtain the value of the parameter $\lambda$ in Eq.(5) by fitting the distribution of stringlet length obtained from simulations. Then, we take the value for the velocity of transverse phonons $v_{T}$ which is obtained by fitting the dispersion relation obtained from simulations. Once these two parameters are known, and fixed by the simulations, our theoretical model provides a parameter-free prediction for the BP frequency in both 2D and 3D amorphous solids, Eqs.(11)-(12). We can then directly compare the predicted value for $\omega_{BP}$, which we will denote as $\omega_{BP}^{th}$, with the numbers obtained from the simulation data (as reported in the original references), which we will label as $\omega_{BP}^{\mathrm{sim}}$.
We start by considering the simulation data for 2D and 3D zero temperature glasses presented in Refs. [35; 36]. In Fig.2, we present two explicit examples performed using the data for the 2D power law model (2DPL) in [35] and the 3D inverse power law model (3DIPL) in [36]. By plausibly assuming that the average velocity of the stringlet propagation coincides with the speed of transverse sound, we obtain a prediction for the BP of the 2D glass given by $\omega_{BP}^{th}\approx 1.344$ which has to be compared with the reported simulation value $\omega_{BP}^{\mathrm{sim}}=1.41$. The difference between the two is less than $5\,\%$.
For the 3D system, the agreement is even superior, showing $\approx 1\,\%$ uncertainty. The predictions of the theoretical model and the compared simulation values are reported in Table 1.
Table 1: The theoretical prediction versus the simulation value for the BP frequency using the exponential fit, Eq.(5). The uncertainty $\Upsilon$ is defined as $\Upsilon=2|\omega_{BP}^{th}-\omega_{BP}^{\mathrm{sim}}|/(\omega_{BP}^{th}+\omega_{BP}^{\mathrm{sim}})$. All quantities presented here are in reduced Lennard-Jones units. The values of the BP frequency are those reported in Refs. [36; 35].
$\lambda$ $v_{T}$ $\omega_{BP}^{th}$ $\omega_{BP}^{\mathrm{sim}}$ $\Upsilon$
3DIPL [36] $1.335$ $4.23$ $2.487$ $2.46$ $1\,\%$
2DPL [35] $3.717$ $4.78$ $1.344$ $1.41$ $4.7\,\%$
The estimates from the theoretical model are in remarkable agreement, i.e., agreement to within $5\,\%$ uncertainty, with the value for the BP frequency obtained from simulations and reported in Refs. [36; 35]. It is emphasized that, once the distribution of length is extracted from the simulation data and the value of the speed of transverse sound is taken from the simulation dispersion relation, the prediction of the BP frequency involves no free parameters.
@@
In order to test further our model, we have extracted the simulation data from [35; 36] and compared them directly to the total VDOS with our predictions. In order to do so, we added a Debye contribution to the stringlet DOS:
$g_{\mathrm{total}}(\omega)=g_{\text{Debye}}(\omega)+g_{s}(\omega)=\mathcal{A}\,\omega^{2}+\mathcal{N}_{s}\,\frac{\,e^{-\frac{\pi v}{\lambda\,\omega}}}{\omega^{2}},$ (19)
where the Debye level $\mathcal{A}$ is fixed using the numerical data, and where we use the normalization constant $\mathcal{N}_{s}$ in the stringlet density of states, Eq.(8), as a fitting parameter. Additionally, as evident from the figures in the bottom panel of Fig.2, a more careful analysis reveals that the fit to an exponential becomes rather uncertain for longer strings which are relatively rare in the simulation sampling, i.e., $l>15$ in reduced LJ units. This was already observed in [59], where a combination of a power-law with an exponential was considered and shown to fit the data better. In order to investigate this point further, we have performed a second analysis in which the stringlet length distribution $p(l)$ is not fitted with the exponential form, Eq.(5), but rather with a power-law corrected expression as in Eq.(13).
We compare our analytic expression for the full reduced density of sates to our analytic model prediction in the top panels of Fig.2. The simulation data are shown with black symbols, the theoretical prediction based on the exponential distribution, Eq.(5), is presented with solid black lines, the theoretical prediction based on the power-law corrected distribution Eq.(13) by the dashed black lines, and finally the gray vertical bar indicates the position of the BP. In the two-dimensional model (2DPL) the comparison between the simulation data and our theoretical prediction is rather good, specially for the exponential fit. The fitting corrected by a power-law term seems to give a slightly better result for frequencies below the BP (corresponding to large $l$ in the distribution function $p(l)$), but a worse result for frequencies around and above the BP. In the three dimensional system, the comparison between the model and the simulations is good below $\omega\approx 4$ (in reduced units), but it becomes poor for large frequencies, much above the BP scale. In particular, the simulation data seem to decay much faster than our theoretical expectations. This discrepancy can be easily understood by noticing that our theoretical prediction does not include any UV cutoff to take into account the microscopic scale of the system. By adding a phenomenological Debye cutoff to the total density of states, $\exp\left(-\omega^{2}/\omega_{D}^{2}\right)$, this discrepancy immediately vanishes as shown by the red dashed curve in the top right panel of Fig.2. Finally, we notice that this discrepancy does not appear in the 2D model because the comparison does not extend to large enough frequencies.
@@ III Damped dynamics in cooled liquids above the glass transition
Inspired by these positive outcomes in the comparisons of the stringlet model to simulations deep in the glass state, we next consider simulations for a simulated Al-Sm metallic glass-forming material in a temperature range above $T_{g}$ discussed previously in [32] (see also [81]). By using the simulation values reported in Table 2, see also Fig.3, we are again able to provide a parameter-free theoretical prediction for the BP frequency in the limit of zero damping, which is shown in the last column of the same table.
Table 2: The parameters for the Al-Sm system at finite temperature [32]. $\lambda$ is reported in nanometers, the velocity in m/s and the frequencies in meV. The last column indicates the theoretical prediction for zero stringlet damping which has to be compared with the previous column reporting the corresponding simulation values. The values of the BP frequency correspond to those reported in Ref.[32].
Al-Sm [32] $\lambda$ $v_{T}$ $\omega_{BP}^{\mathrm{sim}}$ $\omega_{BP}^{th}$ $(\gamma=0)$
@@
The stringlet length distribution and the parameter $\lambda=\langle l\rangle-1$ are shown for different temperatures in the left panel of Fig.3. Moreover, we obtained the values for the density $\rho$ and the shear modulus $G$ from [32; 81] and estimated the transverse speed of sound using the well-known relation:
@@
Figure 3: Left panel: The distribution of the stringlet lengths for the finite temperature Al-Sm system in [32] and the fits to the exponential form, Eq.(5). The inset shows the dependence of the parameter $\lambda$ as a function of temperature. Right panel: The speed of transverse sound, extracted from the simulation data using Eq.(20), as a function of temperature. The values of $\lambda$ and $v_{T}$ are reported explicitly in Table 2.
Figure 4: Left panel: The phenomenological damping parameter used in the theoretical predictions at high temperature, above $T_{g}$, for the Al-Sm system [32]. The colors match those used in the bottom panel. Right panel: The comparison between the theoretical predictions and the simulation results for the finite temperature Al-Sm system in [32]. The black symbols are the simulation data. The orange line indicates the theoretical results in absence of stringlet damping, $\gamma=0$. The pink and red line indicate the theoretical results in presence of damping, as explained in the text.
@@
One qualitative finding is evident from Table 2 (see also Fig.4). Simulation observations cannot be described by a theory neglecting stringlet damping. The theoretical predictions neglecting damping overestimate the observed values of $\omega_{BP}$ by more than a factor of two. The reason for this deviation is easy to understand. As opposed to the simulations of Hu and Tanaka [35; 36], which were performed in the limit of $T=0$ where stringlet damping can be reasonably neglected, the simulations for the Al-Sm metallic glass forming material [32] were performed at relatively high temperatures corresponding to a highly cooled glass-forming metallic glass liquid [32]. At such temperatures, it is unreasonable to assume $\gamma=0$ even as a rough approximation. The treatment of the boson peak at higher temperatures is complicated by the fact that the stringlet length, the transverse sound velocity and the extent of damping described by $\gamma$ can all be expected to vary appreciably with temperature (see Figs. 3 and 4 for the temperature variation observed in the Al-Sm metallic glass-forming liquid). Nevertheless, as we will see below, we can still gain important physical insights into the variation of the density of states and the boson peak with temperature by considering the stringlet model with finite damping $\gamma$ and by considering a simplified description for it.
Before continuing, let us emphasize that our analysis does not take into account the existence of hydrodynamically-defined “defects”, such as the often discussed four-leaf quadrupole defects [82]. Nevertheless, as argued in Ref.[35], these relatively large scale structures apparently correspond to a frequency range much below the boson peak, and thus not germane to the present analysis. In particular, we do not expect this type of defects to have a significant impact in the frequency range near where the boson peak is exhibited in the reduced density of states. It is nevertheless plausible that these additional defects might contribute to the stringlet damping, especially at elevated temperatures and/or large frequencies, and it is also highly likely that these defects are highly significant in connection with understanding plastic deformation in glassy materials. We certainly do not mean to imply these defects are not important in glass-forming materials. Finally, we note that the existence of quadrupole defects defined in relation to elastic distortions in the continuum field describing the material provides a well defined example of how defects might arise without any crystalline structure necessarily existing in the material in order to define a defect. This would seem to imply that a similar non-crystalline based definition might apply to stringlets so that these structures might genuinely be termed “defects”. This possibility remains to be established.
@@
$\frac{g_{s}(\omega)}{\omega^{2}}=\frac{\mathcal{N}_{s}}{\gamma^{2}}\frac{1}{\omega}+\dots$ (21)
@@
Note also that Eq.(21) implies that the VDOS is linear in frequency at low energy in the high temperature liquid state so that our stringlet model exhibits the correct physical behavior in both the low temperature solid and the high temperature liquid states. This behavior is consistent with the experimental observations in classical liquids [83; 84; 85] where it has been rationalized based on the concept of unstable modes [86; 87]. Interestingly, string-like excitations have been also discussed in glass-forming liquids at low temperatures [80; 88; 59], so that it would be important to study further the BP in relation to stringlet dynamics by heating such cooled liquids to watch how the boson peak evolves in shape and eventually disappears. Notably, this disappearance of the boson peak upon heating is well-established experimentally [60; 89; 90; 91; 72; 92]. Similarities to the occurrence of glassy anomalies in materials undergoing commensurate-incommensurate transitions [64; 93; 94; 68; 95] and we suggest that the pursuit of this “analogy” might offer potential insights into the boson peak in glass-forming liquids. We next tentatively explore the extension to higher temperatures where we may expect damping of the stringlets.
@@
We start our discussion about the predictions of our stringlet model at elevated temperatures by noting that an increase of the average stringlet length $\langle l\rangle$ with temperature is not sufficient to reproduce the complete phenomenology of the “softening” of the BP, and the simultaneous emergence of a linear in $\omega$ VDOS as appropriate for liquids [84]. At the outset, we admit that the observed softening of the BP upon heating might involve multiple mechanisms, among which the growth of the average stringlet length $\langle l\rangle=\lambda+1$, the decrease of the velocity $v$, or the increase of the stringlet damping $\gamma$. From the simulation data, only the variation of $\lambda$ and $v$ with $T$ is known, but we cannot currently directly access information about the dependence of $\gamma$ on $T$. Thus, we cannot fully resolve how each mechanism contributes to the BP within the frame of our model without introducing a model of $\gamma(T)$. We introduce a tentative model of $\gamma(T)$ below to address this problem.
@@
Figure 5: Left panel: The stringlet vibrational density of states normalized by Debye law $\omega^{2}$ from Eq.(7). Here, we have fixed $\lambda=1$, $\mathcal{N}_{s}=2\pi$ and we move the damping $\gamma\in[10^{-4},0.25]$ from black to red. The arrows indicate the location of the BP and of the low-frequency $1/\omega$ divergence discussed in Eq.(21). Right panel: The apparent power law emerging at low frequency in presence of damping. In this figure, we have fixed $\lambda=1$, $\mathcal{N}_{s}=2\pi$ and dialed the damping $\gamma\in[0,1]$. A Debye term $g_{\mathrm{Debye}}(\omega)=\omega^{2}$ has been added on top of the stringlet contribution, Eq.(7).